OpenStudy (anonymous):

An office has four copying machines, and the random variable X measures how many of them are in use at a particular moment in time. Suppose that P (X = 0) = 0:08, P (X = 1) = 0:11, P (X = 2) = 0:27, and P (X = 3) = 0:33. What is P (X=4)? Find F(3) Find E(X) Calculate Var(X)

OpenStudy (anonymous):

P (X=4) = 1-(P(x=0)+P(x=1)+P(x=2)+P(x=3)) 1-(.08+.11+.27+.33) P (X=4) = .21